OpenStudy (anonymous):

what is square root of 2 over square root of 3

OpenStudy (anonymous):

\[\sqrt{\left(\begin{matrix}2 \\3 ?\end{matrix}\right)}\]

OpenStudy (anonymous):

sqrt(2/3)

OpenStudy (anonymous):

\[\frac{\sqrt2}{\sqrt3}=\frac{\sqrt2\cdot\sqrt3}{\sqrt3\cdot\sqrt3}=\frac{\sqrt6}{3}\]

OpenStudy (anonymous):

\[\sqrt{2/3}\]